OpenStudy (sleepyjess):

*Calc* Probably making this question harder than it needs to be... question below

OpenStudy (sleepyjess):

Find dx/dt when x = 3, y = 4 and dy/dt = 2 x^2 + y^2 = 25

OpenStudy (sleepyjess):

Derivative: 2x + 2y\(\dfrac{dy}{dt}\) = 0 Do I just need to substitute the values in now or is there something more?

OpenStudy (mayankdevnani):

it should be like this :- \[\large \bf 2x \frac{dx}{dt}+2y \frac{dy}{dt}=0\] plug x = 3, y = 4 and dy/dt = 2 into it and solve for dx/dt

OpenStudy (sleepyjess):

\(2(3)\dfrac{dx}{dt} + 2(4)(2) = 0\) \(6\dfrac{dx}{dt} + 16 = 0 \\ \dfrac{dx}{dt} = 6/16\)

OpenStudy (sleepyjess):

Wait, -16/6

OpenStudy (mayankdevnani):

*-8/3 xD

OpenStudy (sleepyjess):

haha, thanks :)

OpenStudy (mayankdevnani):

anytime :)

OpenStudy (sleepyjess):

Can you help on some related rates problems?

OpenStudy (mayankdevnani):

i'll try